# Split Gibbs Discrete Diffusion Posterior Sampling

### v1

@@ 3.1 Split Gibbs for plug-and-play posterior sampling
$\pi({\mathbf{x}},{\mathbf{z}};\rho)\ \propto\ \exp(-f({\mathbf{z}};{\mathbf{y}})-g({\mathbf{x}})-D({\mathbf{x}},{\mathbf{z}};\rho)),$ (11)
where $D({\mathbf{x}},{\mathbf{z}};\rho)$ measures the distance between ${\mathbf{x}}$ and ${\mathbf{z}}$, and $\rho>0$ is a parameter that controls the strength of regularization. As $\rho\to 0$, $D({\mathbf{x}},{\mathbf{z}};\rho)\to\infty$ for any ${\mathbf{x}}\neq{\mathbf{z}}$, ensuring that both marginal distributions,
$\pi^{X}({\mathbf{x}};\rho):=\int\pi({\mathbf{x}},{\mathbf{z}};\rho)\mathrm{d}{\mathbf{z}},\ \pi^{Z}({\mathbf{z}};\rho):=\int\pi({\mathbf{x}};{\mathbf{z}};\rho)\mathrm{d}{\mathbf{x}},$ (12)
converges to the posterior distribution $p({\mathbf{x}}|{\mathbf{y}})$. The decoupling of the prior and likelihood in Equation 11 enables split Gibbs sampling, which alternates between two steps:
@@ 1. Likelihood sampling step:
${\mathbf{z}}^{(k)}\sim\pi({\mathbf{x}}={\mathbf{x}}^{(k)},{\mathbf{z}};\rho)\ \propto\ \exp(-f({\mathbf{z}};{\mathbf{y}})-D({\mathbf{x}}^{(k)},{\mathbf{z}};\rho)),$ (13)
@@ 2. Prior sampling step:
${\mathbf{x}}^{(k+1)}\sim\pi({\mathbf{x}},{\mathbf{z}}={\mathbf{z}}^{(k)};\rho)\ \propto\ \exp(-g({\mathbf{x}})-D({\mathbf{x}},{\mathbf{z}}^{(k)};\rho)).$ (14)
@@
Prior work has studied split Gibbs posterior sampling for continuous diffusion models (Coeurdoux et al., 2023; Wu et al., 2024; Xu & Chi, 2024). These methods specify the regularization potential as $D({\mathbf{x}},{\mathbf{z}};\rho)=-\frac{\|{\mathbf{x}}-{\mathbf{z}}\|_{2}^{2}}{2\rho^{2}}$, which transforms Equation 14 into a Gaussian denoising problem solvable by a diffusion model.
In the following sections, we explore how a specific choice of the potential $D({\mathbf{x}},{\mathbf{z}};\rho)$ links split Gibbs sampling to discrete diffusion models and discuss how the prior and likelihood sampling steps work in discrete-state spaces.
@@ 3.2 Prior Step with Discrete Diffusion Models
Suppose $p({\mathbf{x}})$ is a discrete-state distribution over $\mathcal{X}^{D}$ modeled by a diffusion process. We consider a discrete diffusion model with uniform transition kernel ${\bm{Q}}_{t}=\frac{1}{N}\mathbf{1}\mathbf{1}^{T}-{\bm{I}}$. We specify the potential function $D({\mathbf{x}};{\mathbf{z}};\rho)$ as:
$D({\mathbf{x}},{\mathbf{z}};\rho):=d({\mathbf{x}},{\mathbf{z}})\log\frac{1+(N-1)e^{-\rho}}{(N-1)(1-e^{-\rho})},$ (15)
where $d({\mathbf{x}},{\mathbf{z}})$ denotes the Hamming distance between ${\mathbf{x}}$ and ${\mathbf{z}}$. When $\rho\to 0^{+}$, the regularization potential $D({\mathbf{x}},{\mathbf{z}};\rho)$ goes to infinity unless $d({\mathbf{x}},{\mathbf{z}})=0$, ensuring the convergence of marginal distributions to $p({\mathbf{x}}|{\mathbf{y}})$. Given Equation 15, the prior step from Equation 14 is:
${\mathbf{x}}^{(k+1)}$ $\sim\pi({\mathbf{x}},{\mathbf{z}}={\mathbf{z}}^{(k)};\rho)$
$\ \propto\ p_{0}({\mathbf{x}})\left(\frac{\tilde{\beta}}{1-\tilde{\beta}}\right)^{d({\mathbf{z}}^{(k)},{\mathbf{x}})},$ (16)
where $\tilde{\beta}=\frac{N-1}{N}(1-e^{-\rho})$. On the other hand, the distribution of clean data ${\mathbf{x}}_{0}$ conditioned on ${\mathbf{x}}_{t}$ is given by:
$p({\mathbf{x}}_{0}|{\mathbf{x}}_{t})$ $\ \propto\ p_{0}({\mathbf{x}}_{0})p({\mathbf{x}}_{t}|{\mathbf{x}}_{0})$
$\ \propto\ p_{0}({\mathbf{x}}_{0})\beta_{t}^{d({\mathbf{x}}_{t},{\mathbf{x}}_{0})}(1-\beta_{t})^{D-d({\mathbf{x}}_{t},{\mathbf{x}}_{0})},$ (17)
where $\beta_{t}=\frac{N-1}{N}(1-e^{-\sigma_{t}})$. Thus, sampling from Equation 16 is equivalent to unconditional generation from $p({\mathbf{x}}_{0}|{\mathbf{x}}_{t})$ when $\tilde{\beta}=\beta_{t}$, i.e., $\rho=\sigma_{t}$. Therefore, we can solve the prior sampling problem by simulating a partial discrete diffusion sampler starting from $\sigma_{t}=\rho$ and ${\mathbf{x}}_{t}={\mathbf{z}}^{(k)}$.
@@ 3.3 Likelihood Sampling Step
With the potential function $D({\mathbf{x}},{\mathbf{z}};\rho)$ specified in Section 3.2, at iteration $k$, the likelihood sampling step from Equation 13 can be written as:
${\mathbf{z}}^{(k)}$ $\sim\pi({\mathbf{x}}={\mathbf{x}}^{(k)},{\mathbf{z}};\rho)$
$\ \propto\ \exp\left(-f({\mathbf{z}};{\mathbf{y}})-d({\mathbf{x}},{\mathbf{z}})\log\frac{1+(N-1)e^{-\rho}}{(N-1)(1-e^{-\rho})}\right).$ (18)
Since the unnormalized probability density function of $\pi({\mathbf{x}}={\mathbf{x}}^{(k)},{\mathbf{z}};\rho)$ is accessible, we can efficiently sample from Equation 18 using Markov Chain Monte Carlo methods. Specifically, we initialize ${\mathbf{z}}^{(k)}_{0}={\mathbf{x}}^{(k)}$ and run the Metropolis-Hasting algorithm to sample from Equation 18. The updating rule is given by:
${\mathbf{z}}^{(k)}_{j+1}=\begin{cases}\tilde{\mathbf{z}}_{j},\quad\ \ \,\mathrm{w.p.}\ \alpha({\mathbf{z}}^{(k)}_{j}\to\tilde{\mathbf{z}}_{j}),\\ {\mathbf{z}}^{(k)}_{j},\quad\mathrm{otherwise},\end{cases}$ (19)
where $\tilde{\mathbf{z}}_{j}$ is a proposal obtained by randomly flipping $s$ tokens in ${\mathbf{z}}^{(k)}_{j}$, and $\alpha({\mathbf{z}}^{(k)}_{j}\to\tilde{\mathbf{z}}_{j})$ denotes the acceptance probability of the proposal, given by:
$\alpha({\mathbf{z}}^{(k)}_{j}\to\tilde{\mathbf{z}}_{j})=\min\left(1,\frac{\pi({\mathbf{x}}={\mathbf{x}}^{(k)},\tilde{\mathbf{z}}_{j};\rho)}{\pi({\mathbf{x}}={\mathbf{x}}^{(k)},{\mathbf{z}}^{(k)}_{j};\rho)}\right).$ (20)
We repeat the Metropolis-Hastings update in Equation 19 for $T$ steps and set ${\mathbf{z}}^{(k)}={\mathbf{z}}^{(k)}_{T}$.
Algorithm 1 Split Gibbs Diffusion Posterior Sampling
  Require: Concrete score model ${\mathbf{s}}_{\theta}$, measurement ${\mathbf{y}}$, noise schedule $\{\rho_{k}\}_{k=0}^{K-1}$.
@@
     Likelihood step following Equation 18
${\mathbf{z}}^{(k)}\sim\pi({\mathbf{x}}={\mathbf{x}}^{(k)},{\mathbf{z}};\rho_{k}).$
     Prior step following Equation 16.
${\mathbf{x}}^{(k+1)}\sim\pi({\mathbf{x}},{\mathbf{z}}={\mathbf{z}}^{(k)};\rho_{k})$
@@
SG-DPS 0.149 0.125 0.214 0.222 0.334 0.365
Table 1: Quantitative evaluation of posterior sampling on synthetic dataset for sequence lengths of $D=2,5$ and $10$. The Hellinger distance and total variance are computed between the empirical distribution of 10k samples and the true posterior distribution on the first two dimensions.
3.4 Overall Algorithm
We now summarize the complete SG-DPS algorithm. Like in Wu et al. (2024), our algorithm alternates between likelihood steps and prior steps while employing an annealing schedule $\{\rho_{k}\}$, which starts at a large $\rho_{0}$ and gradually decays to $\rho_{K}\to 0$. This annealing scheme accelerates the mixing time of the Markov chain and ensures the convergence of $\pi^{X}({\mathbf{x}};\rho)$ and $\pi^{Z}({\mathbf{z}};\rho)$ to $p({\mathbf{x}}|{\mathbf{y}})$ as $\rho\to 0$. Unlike prior work that uses uncontrolled approximations such as surrogate value functions,  SG-DPS does not rely on any approximation and eventually converges to the true posterior distribution. Consequently, SG-DPS can be viewed as a principled approach for posterior sampling using discrete diffusion models, leading to empirical benefits as shown in Section 4. We present the complete pseudocode of our method in Algorithm 1 and further explain the choice of the noise schedule in Section 4.6.
4 Experiments
We conduct experiments to demonstrate the effectiveness of our algorithm on various posterior sampling tasks in discrete-state spaces, including discrete inverse problems and reward/loss-guided generation.
4.1 Experimental Setup
We evaluate our method across multiple domains, including synthetic data, discrete image prior, monophonic music, and DNA sequences.
Pretrained models. For synthetic data, we use a closed-form concrete score function. For real datasets, we train a SEDD (Lou et al., 2024) model with the uniform transition kernel. These SEDD models are trained using AdamW (Loshchilov & Hutter, 2019) optimizer with a batch size of $32$ and a learning rate of $3\times 10^{-4}$.
Baselines. We compare our methods to existing approaches for discrete diffusion posterior sampling: DPS (Chung et al., 2023), SVDD (Li et al., 2024). SVDD is originally designed to sample from a reward-guided distribution $p^{\alpha}({\mathbf{x}})\ \propto\ p({\mathbf{x}})\exp(r({\mathbf{x}})/\alpha)$. To adapt it for posterior sampling, we let $r({\mathbf{x}})=\log p({\mathbf{y}}|{\mathbf{x}})$ and use a temperature of $\alpha=1$, so that $p^{\alpha}({\mathbf{x}})=p({\mathbf{x}}|{\mathbf{y}})$. DPS relies on gradient back-propagation through both the likelihood function and the pretrained diffusion model, making it unsuitable for discrete-domain posterior sampling. To adapt DPS for discrete diffusion models, we replace the gradient term $\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{y}}|{\mathbf{x}}_{t})$ with a guidance rate matrix ${\bm{R}}_{t}^{\mathbf{y}}=[\frac{p_{t}({\mathbf{y}}|\tilde{\mathbf{x}})}{p_{t}({\mathbf{y}}|{\mathbf{x}})}]_{{\mathbf{x}}\tilde{\mathbf{x}}}$, which is effectively a finite-difference approximation of a “discrete derivative”. Implementation details are provided in Section A.2.
SG-DPS implementation details. We implement our method with a total of $K=100$ iterations. In each prior sampling step, we simulate the reverse continuous-time Markov chain with $L=20$ steps. The likelihood function for inverse problems is defined as
$p({\mathbf{y}}|{\mathbf{x}})\ \propto\ \exp(-\|{\bm{G}}({\mathbf{x}})-{\mathbf{y}}\|_{0}/\sigma_{{\mathbf{y}}}).$ (21)
We set $\sigma_{{\mathbf{y}}}=0.1$ for most experiments. Additional details on hyperparameter choices are provided in Section A.3.

### v2

@@ 3.1 Split Gibbs for plug-and-play posterior sampling
$\pi({\mathbf{x}},{\mathbf{z}};\eta)\ \propto\ \exp(-f({\mathbf{z}};{\mathbf{y}})-g({\mathbf{x}})-D({\mathbf{x}},{\mathbf{z}};\eta)),$ (7)
where $D({\mathbf{x}},{\mathbf{z}};\eta)$ measures the distance between ${\mathbf{x}}$ and ${\mathbf{z}}$, and $\eta>0$ is a parameter that controls the strength of regularization. While prior works [8, 48, 49] consider $D({\mathbf{x}},{\mathbf{z}};\eta)=\frac{\|{\mathbf{x}}-{\mathbf{z}}\|^{2}_{2}}{2\eta^{2}}$, the potential function can be generalized to any continuous function $D$, such that $D({\mathbf{x}},{\mathbf{z}};\eta)\to\infty$ as $\eta\to 0$ for any ${\mathbf{x}}\neq{\mathbf{z}}$. As shown in Section A.3, this ensures that both marginal distributions,
$\pi^{X}({\mathbf{x}};\eta):=\int\pi({\mathbf{x}},{\mathbf{z}};\eta)\mathrm{d}{\mathbf{z}},\ \ \text{and}\ \ \pi^{Z}({\mathbf{z}};\eta):=\int\pi({\mathbf{x}};{\mathbf{z}};\eta)\mathrm{d}{\mathbf{x}}$
converges to the posterior distribution $p({\mathbf{x}}|{\mathbf{y}})$. The decoupling of the prior and likelihood in Eq. 7 enables split Gibbs sampling, which alternates between two steps:
@@ 1. Likelihood sampling step:
${\mathbf{z}}^{(k)}\sim\pi({\mathbf{x}}={\mathbf{x}}^{(k)},{\mathbf{z}};\eta)\ \propto\ \exp(-f({\mathbf{z}};{\mathbf{y}})-D({\mathbf{x}}^{(k)},{\mathbf{z}};\eta)),$ (8)
@@ 2. Prior sampling step:
${\mathbf{x}}^{(k+1)}\sim\pi({\mathbf{x}},{\mathbf{z}}={\mathbf{z}}^{(k)};\eta)\ \propto\ \exp(-g({\mathbf{x}})-D({\mathbf{x}},{\mathbf{z}}^{(k)};\eta)).$ (9)
@@
Prior work has studied split Gibbs posterior sampling for continuous diffusion models [8, 48, 49]. These methods specify the regularization potential as $D({\mathbf{x}},{\mathbf{z}};\eta)=\frac{\|{\mathbf{x}}-{\mathbf{z}}\|_{2}^{2}}{2\eta^{2}}$, which transforms Eq. 9 into a Gaussian denoising problem solvable by a diffusion model. A key challenge, which we tackle, is developing a performant approach for the discrete diffusion setting that is easy to implement, and enjoys rigorous guarantees (i.e., sampling from the correct posterior distribution).
@@ 3.2 Prior Step with Discrete Diffusion Models
Suppose $p({\mathbf{x}})$ is a discrete-state distribution over $\mathcal{X}^{D}$ modeled by a diffusion process. We consider a discrete diffusion model with uniform transition kernel ${\bm{Q}}_{t}^{\mathsf{fw}}=\frac{1}{N}\mathbf{1}\mathbf{1}^{T}-{\bm{I}}$. To connect split Gibbs sampling to discrete diffusion models, we specify the potential function $D({\mathbf{x}};{\mathbf{z}};\eta)$ as:
$D({\mathbf{x}},{\mathbf{z}};\eta):=d({\mathbf{x}},{\mathbf{z}})\log\frac{1+(N-1)e^{-\eta}}{(N-1)(1-e^{-\eta})}$ (10)
where $d({\mathbf{x}},{\mathbf{z}})$ denotes the Hamming distance between ${\mathbf{x}}$ and ${\mathbf{z}}$. When $\eta\to 0^{+}$, the regularization potential $D({\mathbf{x}},{\mathbf{z}};\eta)$ goes to infinity unless $d({\mathbf{x}},{\mathbf{z}})=0$, ensuring the convergence of marginal distributions to $p({\mathbf{x}}|{\mathbf{y}})$. Given Eq. 10, the prior step from Eq. 9 can be written as
${\mathbf{x}}^{(k+1)}$ $\sim\pi({\mathbf{x}},{\mathbf{z}}={\mathbf{z}}^{(k)};\eta)\ \propto\ p_{0}({\mathbf{x}})\left(\tilde{\beta}/(1-\tilde{\beta})\right)^{d({\mathbf{z}}^{(k)},{\mathbf{x}})}$ (11)
where $\tilde{\beta}=\frac{N-1}{N}(1-e^{-\eta})$. On the other hand, the distribution of clean data ${\mathbf{x}}_{0}$ conditioned on ${\mathbf{x}}_{t}$ is given by:
$p({\mathbf{x}}_{0}|{\mathbf{x}}_{t})$ $\ \propto\ p_{0}({\mathbf{x}}_{0})p({\mathbf{x}}_{t}|{\mathbf{x}}_{0})\ \propto\ p_{0}({\mathbf{x}}_{0})\beta_{t}^{d({\mathbf{x}}_{t},{\mathbf{x}}_{0})}(1-\beta_{t})^{D-d({\mathbf{x}}_{t},{\mathbf{x}}_{0})}$ (12)
where $\beta_{t}=\frac{N-1}{N}(1-e^{-\sigma_{t}})$. Thus, sampling from Eq. 11 is equivalent to unconditional generation from $p({\mathbf{x}}_{0}|{\mathbf{x}}_{t})$ when $\tilde{\beta}=\beta_{t}$, i.e., $\eta=\sigma_{t}$. Therefore, we can solve the prior sampling problem by simulating a partial discrete diffusion sampler starting from $\sigma_{t}=\eta$ and ${\mathbf{x}}_{t}={\mathbf{z}}^{(k)}$.
@@ 3.3 Likelihood Sampling Step
With the potential function $D({\mathbf{x}},{\mathbf{z}};\eta)$ specified in Section 3.2, at iteration $k$, the likelihood sampling step from Eq. 8 can be written as:
${\mathbf{z}}^{(k)}\sim\pi({\mathbf{x}}={\mathbf{x}}^{(k)},{\mathbf{z}};\eta)\ \propto\ \exp\left(-f({\mathbf{z}};{\mathbf{y}})-d({\mathbf{x}},{\mathbf{z}})\log\frac{1+(N-1)e^{-\eta}}{(N-1)(1-e^{-\eta})}\right).$ (13)
Since the unnormalized probability density function of $\pi({\mathbf{x}}={\mathbf{x}}^{(k)},{\mathbf{z}};\eta)$ is accessible, we can efficiently sample from Eq. 13 using Markov Chain Monte Carlo methods.
Algorithm 1 Split Gibbs Discrete Diffusion Posterior Sampling (SGDD)
  Require: Concrete score model ${\mathbf{s}}_{\theta}$, measurement ${\mathbf{y}}$, noise schedule $\{\eta_{k}\}_{k=0}^{K-1}$.
@@
     Likelihood step following Eq. 13: ${\mathbf{z}}^{(k)}\sim\pi({\mathbf{x}}={\mathbf{x}}^{(k)},{\mathbf{z}};\eta_{k}).$
     Prior step following Eq. 11: ${\mathbf{x}}^{(k+1)}\sim\pi({\mathbf{x}},{\mathbf{z}}={\mathbf{z}}^{(k)};\eta_{k})$
@@
3.4 Overall Algorithm
We now summarize the complete SGDD algorithm. Like in Wu et al. [48], our algorithm alternates between likelihood steps and prior steps while employing an annealing schedule $\{\eta_{k}\}$, which starts at a large $\eta_{0}$ and gradually decays to $\eta_{K}\to 0$. This annealing scheme accelerates the mixing time of the Markov chain, which we further provide some theoretical intuition in Section B.1; and ensures the convergence of $\pi^{X}({\mathbf{x}};\eta)$ and $\pi^{Z}({\mathbf{z}};\eta)$ to $p({\mathbf{x}}|{\mathbf{y}})$ as $\eta\to 0$. We present the complete pseudocode of our method in Algorithm 1 and further explain the choice of the noise schedule in Section B.1.
3.5 Convergence of Split Gibbs Diffusion Posterior Sampling
We provide theoretical guarantees on the convergence of SGDD. For two probability measures in a finite $\mathcal{X}$, we consider the Kullback-Leibler (KL) divergence and the relative Fisher information (a.k.a. Fisher divergence), respectively, as (where we define $f:=\mu/\pi$)
$\mathsf{KL}(\mu\|\pi):=\mathbb{E}_{{\mathbf{x}}\sim\mu}[\ln\frac{\mu}{\pi}({\mathbf{x}})],\ \mathsf{FI}_{{\bm{Q}}}(\mu\|\pi):=\sum_{{\mathbf{x}}_{i},{\mathbf{x}}_{j}\in\mathcal{X}}\pi({\mathbf{x}}_{i}){\bm{Q}}^{[j,i]}\left(f({\mathbf{x}}_{j})-f({\mathbf{x}}_{i})-f({\mathbf{x}}_{i})\log\frac{f({\mathbf{x}}_{j})}{f({\mathbf{x}}_{i})}\right).$
Both divergences are nonnegative and equal to zero if and only if $\mu=\pi$, when ${\bm{Q}}$ is irreducible. Fisher information has been widely used to analyze the convergence of sampling algorithms in continuous spaces due to its inherent connection to KL divergence  [2, 41, 48]. In the continuous case, the Fisher information can be written as a quadratic form $\mathsf{FI}(\mu\|\pi)=\mathbb{E}_{\mu}\|\nabla\log(\mu/\pi)\|^{2}$. However, it does not have an analogous form in finite spaces, which poses additional challenges to the analysis of SGDD. To address this challenge, we adopt a generalized definition of Fisher information [3, 17] and analyze the convergence of SGDD to the stationary distribution using a more general technique that encompasses both continuous and discrete settings.
We define a distribution $\mu_{t}$ over $\mathcal{X}$ that evolves according to likelihood steps and prior steps alternatively. We assume each likelihood step is implemented with the Metropolis-Hastings algorithm, and that each prior step is solved by the Euler method with an approximated score function. We compare $\mu_{t}$ to the continuous-time stationary distribution $\pi_{t}$, which alternates between $\pi^{X}$ and $\pi^{Z}$.
Theorem 1.
Consider running $K$ iterations of SGDD with a fixed $\eta>0$ and an estimated concrete score $s_{\theta}({\mathbf{x}};t)$, and suppose that each prior step is solved by an $H$ step Euler method. Let $t^{*}>0$ with $\sigma(t^{*})=\eta$. Define $\pi_{t}$ and $\mu_{t}$ as stationary and non-stationary distributions. Over $K$ iterations of SGDD, the average relative Fisher information between $\mu_{t}$ and $\pi_{t}$ has
$\frac{1}{K}\sum_{k=0}^{K-1}\underbrace{\frac{1}{t^{*}}\int_{k(t^{*}+1)+1}^{(k+1)(t^{*}+1)}\mathsf{FI}_{{\bm{Q}}_{t}}(\mu_{t}\|\pi_{t})}_{\textit{average Fisher divergence in the k-th prior step}}\leq\underbrace{\frac{2\mathsf{KL}(\mu_{0}\|\pi_{0})}{Kt^{*}}}_{\textit{convergence by}\ O(1/K)}+\underbrace{\frac{4M\epsilon}{c}}_{\textit{score error}}+\underbrace{\frac{2MLt^{*}}{cH}}_{\textit{discretization error}}.$ (14)
where $\|\frac{s_{\theta}(\cdot,t)-s(\cdot,t)}{s(\cdot,t)}\|_{\infty}\leq\epsilon<1$, and $L,M,c$ are positive constants defined in Appendix A.
We include the proof in Appendix A. Theorem 1 states that the average Fisher divergence of the non-stationary process to the stationary process in all prior steps converges at the rate of $O(1/K)$, up to a constant error term. This result extends our theoretical understanding of diffusion posterior sampling by generalizing the analysis [41, 48] on SDEs to general Markov processes using the free-energy-rate-functional-relative-Fisher (FIR) inequality [16]. Moreover, compared to existing analyses for continuous diffusion models, our analysis accounts not only for the imperfect score function but also for the discretization error in solving CTMCs as well.
Unlike SMC-based methods that use uncontrolled assumptions such as surrogate value functions and an infinite number of particles, SGDD does not rely on such approximations and eventually converges to the posterior distribution with robustness to score approximation and discretization error. Consequently, SGDD can be viewed as a natural and principled approach for posterior sampling using discrete diffusion models, leading to empirical benefits as we demonstrate in Section 4.
4 Experiments
Table 1: A summary of posterior sampling tasks considered in our experiments.
Domain Synthetic DNA enhancers Images Monophonic music
Task type Inverse problem Reward guidance Inverse problem Inverse problem
Dimensionality $50^{10}$ $4^{200}$ $2^{1024}$ $129^{256}$
We conduct experiments to demonstrate the effectiveness of our algorithm on various posterior sampling tasks in discrete-state spaces, including discrete inverse problems and conditional generation guided by a reward function.
4.1 Experimental Setup
We evaluate our method across multiple domains, including synthetic data, DNA sequences, discrete image prior, and monophonic music, as listed in Table 1. We define the likelihood function for all inverse problems as $p({\mathbf{y}}|{\mathbf{x}})\ \propto\ \exp(-\|{\bm{G}}({\mathbf{x}})-{\mathbf{y}}\|/\sigma_{\mathbf{y}})$, while the likelihood functions for reward guiding tasks are $p({\mathbf{y}}|{\mathbf{x}})\ \propto\ \exp(\beta r({\mathbf{x}}))$.
Pretrained models. For synthetic data, we use a closed-form concrete score function. For real datasets, we train a SEDD [29] model with the uniform transition kernel (details in Appendix C).
Baselines. We compare our methods to existing approaches for discrete diffusion posterior sampling: DPS [7], SVDD-PM [27], and SMC [47]. The original version of DPS relies on gradient back-propagation through both the likelihood function and the pretrained diffusion model, making it unsuitable for discrete-domain posterior sampling. In our experiments, we consider its discrete analogy by replacing the gradient with a guidance rate matrix. SVDD-PM is originally designed to sample from a reward-guided distribution $p^{\beta}({\mathbf{x}})\ \propto\ p({\mathbf{x}})\exp(\beta r({\mathbf{x}}))$, but it is also adaptable to solving inverse problems. More implementation details are elaborated in Section C.2.
SGDD implementation details. We implement our method with a total of $K$ iterations. In each prior sampling step, we simulate the reverse continuous-time Markov chain with $H=20$ steps. Additional details on hyperparameter choices are provided in Section C.3.
Table 2: Quantitative evaluation of posterior sampling on synthetic dataset for sequence lengths of $D=2,5$ and $10$. The Hellinger distance and total variation are computed between the empirical distribution of 10k samples and the true posterior distribution on the first two dimensions.
@@ 4.1 Experimental Setup
SGDD 0.149 0.125 0.214 0.222 0.334 0.365
4.2 Controlled Study using Synthetic Data
To evaluate the accuracy of SGDD in posterior sampling, we first conduct a case study on a synthetic task where the true posterior distribution is accessible. Let ${\mathbf{x}}\in\mathcal{X}^{D}$ follow a prior distribution $p({\mathbf{x}})$, obtained by discretizing a Gaussian distribution $\mathcal{N}(0,\sigma^{2}{\bm{I}}_{D})$ over an $D$-dimensional grid. We define a forward model by ${\bm{G}}({\mathbf{x}})=\|f({\mathbf{x}})\|_{1}$ where $f({\mathbf{x}})$ maps ${\mathbf{x}}$ to $\mathbb{R}^{D}$.
We draw 10k independent samples with each algorithm and compute the frequency map for the first two dimensions of ${\mathbf{x}}$. We measure the Hellinger distance and total variation distance between the empirical distribution and the ground truth posterior distribution. As shown in Table 2, our method achieves substantially more accurate posterior sampling compared to baseline methods. We include visualization of the generated samples in Appendix D.3 (Fig. 8) by projecting samples onto the first two dimensions and computing their heatmaps. When the number of dimensions is small, e.g., $D=2$, all methods approximate the true posterior distribution well. However, as the dimensionality increases, DPS, SMC, and SVDD-PM struggle to capture the posterior and tend to degenerate to the prior distribution. While SGDD consistently outperforms baseline methods in terms of Hellinger distance and total variation distance, it also shows a slight deviation from the posterior distribution as the dimensionality increases.
4.3 Guided DNAs Generation
We evaluate our method for generating regulatory DNA sequences that drive gene expression in specific cell types, which is an important task for cell and gene therapy and commonly studied in the literature [27, 42, 45].
We train a discrete diffusion model [29] on a publicly available dataset from Gosai et al. [15] that consists of $\sim$700k DNA sequences. The expression driven by each sequence in human cell lines is measured by massively parallel reporter assays (MPRAs). Following Wang et al. [45], we split the dataset into two subsets and train surrogate reward oracles to predict the activity level in the HepG2 cell line. We use the oracle trained on the training set, $r(\cdot)$, to guide our diffusion model, and evaluate its performance using the other one trained on the test dataset. The likelihood function is therefore defined as $p({\mathbf{y}}|{\mathbf{x}})\ \propto\ \exp(\beta r({\mathbf{x}}))$.
Evaluation metrics. We evaluate the quality of generated samples using the following metrics:
  • Predicted activity (Pred-Activity). We evaluate the activity level of generated sequences using the oracle trained on the test dataset. This corresponds to the reward function that we aim to optimize.
  • Binary classification on chromatin accessibility (ATAC-Acc). Following [45], we use a binary classifier trained on chromatin accessibility data in the HepG2 cell line. As active enhancers should have accessible chromatin, ATAC-Acc can be used to validate the generated sequences.
  • Log-likelihood. We estimate the log-likelihood of the generated samples using the ELBO of a pretrained masked diffusion model [45]. This regulates the generated samples to stay close to the prior distribution of natural DNA sequences.
We include more baseline methods that require additional training for this task, including a discrete diffusion fine-tuning method, DRAKES [45], classifier guidance (CG) for discrete diffusion models [32], and twisted diffusion sampler (TDS) [47] which uses CG as the proposal function for sequential Monte Carlo sampling.
Table 3: Quantitative results for guided DNA (enhancers) sampling. We generate 640 independent samples with different algorithms according to the reward-shaped posterior distribution. We report the mean and median of the enhancer’s activity computed by the surrogate reward.
[…]
SVDD-PM $5.41$ $5.08_{\pm 1.51}$ $49.9$$\textbf{241}_{\pm 47}$
SMC $4.15$ $3.99_{\pm 1.49}$ $39.9$$259_{\pm 13}$
CG $2.90$ $2.76_{\pm 0.94}$ $\ \ 0.0$$265_{\pm 5\hskip 5.0pt}$
TDS $4.64$ $4.62_{\pm 1.09}$ $45.3$$257_{\pm 9\hskip 5.0pt}$
DRAKES w/o KL $6.44$ $6.28_{\pm 0.76}$ $82.5$$281_{\pm 6\hskip 5.0pt}$
DRAKES w/ KL $5.61$ $5.24_{\pm 1.39}$ $92.5$$264_{\pm 9\hskip 5.0pt}$
SGDD $(\beta=30)$ $8.82$ $\underline{8.69}_{\pm 1.12}$ $91.7$$\underline{246}_{\pm 41}$
SGDD $(\beta=50)$ 9.14 $\textbf{8.96}_{\pm 1.17}$ 93.0 -$261_{\pm 29}$
Results. We generate $640$ DNA samples using different algorithms and compute the median and average Pred-Activity of the generated sequences. As shown in Table 3, our method achieves the highest Pred-Activity (reward values) compared to existing baselines, including 42% higher median reward compared to the previous state-of-the-art method.
Moreover, as SGDD samples from the posterior distribution, $p({\mathbf{x}})\exp(\beta r({\mathbf{x}}))/Z$, the guidance strength $\beta$ controls the strength of guidance imposed in the sampling algorithm. We instantiate SGDD with $\beta$ ranging from $3$ to $100$ and calculate the Pred-Activity and Log-likelihood of the generated samples. We plot the distribution of the samples’ Pred-Activity with various $\beta$ in Fig. 3. In general, using a larger $\beta$ results in higher Pred-Activity (reward), at the expense of a lower likelihood as the target distribution deviates further from the prior distribution.
Figure 2: Generating DNA sequences involves trade-offs between targeted objectives and consistency with the prior distribution. SGDD is plotted in blue and darker colors mean higher $\beta$.
Figure 3: Diversified samples when the measurement ${\mathbf{y}}$ is sparse. Samples are generated by SGDD when solving an MNIST inpainting task.
4.4 Discrete Image Inverse Problems
Table 4: Quantitative results for XOR and AND problems on discretized MNIST. We report the mean and standard deviation of PSNR and class accuracy across 1k generated samples. SGDD demonstrates superior performance on both tasks.